OpenStudy (anonymous):

A sample of 3.30 grams of an ideal gas at 150.0 C and 1.25 atmospheres pressure has a volume of 2.00 liters. What is the molar mass of the gas? The gas constant, R, is 0.0821 (L*atm)/(mol*K) A) 0.0218 grams/mole B) 16.2 grams/mole C) 37.0 grams/mole D) 45.8 grams/mole E) 71.6 grams/mole

OpenStudy (matt101):

Remember the gas law, PV=nRT. The question gives you values for P, V, R, and T. The only thing we're missing is n. However, we're also given a mass, m, of the sample, and we want to find molar mass, MM. The nice thing is that n=m/MM. So, we can update the gas law to use m/MM instead of just n: PV=(m/MM)RT. Now you can just plug in all the variables you're given and solve for MM! Let me know what you get!

OpenStudy (anonymous):

I got 45.8 as the answer!

OpenStudy (matt101):

Looks good to me!

OpenStudy (anonymous):

Thank you so much!